BRICS Notes Series, 1994

December 10, 2003

This document is also available as PostScript and DVI.

References

[NS-94-6] Abstract, PostScript.

Uffe H. Engberg, Kim G. Larsen, and Peter D. Mosses, editors. Proceedings of the 6th Nordic Workshop on Programming Theory (Aarhus, Denmark, 17–19 October, 1994), December 1994. v+483pp.

[NS-94-5] Abstract, PDF.

Andrew M. Pitts. Some Notes on Inductive and Co-Inductive Techniques in the Semantics of Functional Programs, DRAFT VERSION. December 1994. vi+135 pp.

[NS-94-4] Abstract, PostScript, DVI.

Peter D. Mosses, editor. Abstracts of the 6th Nordic Workshop on PROGRAMMING THEORY (Aarhus, Denmark, 17–19 October, 1994), October 1994. v+52 pp.

[NS-94-3] Abstract, PDF.

Sven Skyum, editor. Complexity Theory: Present and Future (Aarhus, Denmark, 15–18 August, 1994), September 1994. v+213 pp.

[NS-94-2] Abstract, PDF.

David A. Basin. Induction Based on Rippling and Proof Planning. Mini-Course. August 1994. 62 pp.

[NS-94-1] Abstract, PDF.

Peter D. Mosses, editor. Proc. 1st International Workshop on Action Semantics (Edinburgh, 14 April, 1994), May 1994. 145 pp.